# SIGGesture: Generalized Co-Speech Gesture Synthesis via Semantic Injection with Large-Scale Pre-Training Diffusion Models

### v1

@@ Abstract
The automated synthesis of high-quality 3D gestures from speech is of significant value in virtual humans and gaming. Previous methods focus on synthesizing gestures that are synchronized with speech rhythm, yet they frequently overlook the inclusion of semantic gestures. These are sparse and follow a long-tailed distribution across the gesture sequence, making them difficult to learn in an end-to-end manner. Moreover, generating gestures, rhythmically aligned with speech, faces a significant issue that cannot be generalized to in-the-wild speeches. To address these issues, we introduce SIGGesture, a novel diffusion-based approach for synthesizing realistic gestures that are of both high quality and semantically pertinent. Specifically, we firstly build a strong diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels. Secondly, we leverage the powerful generalization capabilities of Large Language Models (LLMs) to generate proper semantic gestures for the various speech content. Finally, we propose a semantic injection module to infuse semantic information into the synthesized results during diffusion reverse process. Extensive experiments demonstrate that the proposed SIGGesture significantly outperforms existing baselines and shows excellent generalization and controllability.
Keywords:
@@
Co-speech gesture synthesis Semantic gestures Large Language Models Diffusion models
1 Introduction
@@
Body gesture can greatly improve people’s expressiveness in communication, which can emphasize or convey specific semantics with non-verb body movements [8, 18, 49, 36]. Besides, high-quality 3D gesture animation can present vivid virtual characters that help listeners to concentrate and improve the intimacy between humans and virtual characters. Therefore, there is a growing need for automatic high-quality co-speech gesture synthesis in many applications, including Virtual Humans, non-player gaming characters, and robot assistants. Many researchers have proposed lots of approaches including rule-based methods [24, 35] and data-driven methods [52, 53, 2].
@@
Figure 1: The framework of the proposed method. The upper shows the training details of diffusion process and denoising process. The lower shows the inference process of synthesizing the gesture from the conditions. Specifically, given audio feature, speaker identity, and speech content, the audio feature and speaker identity are directly fed into the denoising network. The speech content are as input in the interaction with LLM, and the LLM will generate a set of candidate semantic gestures from the input text for the next semantic injection process.
@@
3D co-speech gesture generation aims to generate gestures with natural movements, accurate rhythm, and clear semantics based on a given speech. Rule-based methods [24, 35] achieve this goal by hard-rules, such as keyword "good" with a simple gesture "thump up". However, these approaches can only deal with simple semantic gestures, which are lack of naturalness, smoothness, and diversity. The principal reason for this is that generating co-speech gestures poses a challenge due to its inherently complex many-to-many mapping. Recently, deep learning with large-scale dataset has achieved many breakthroughs, especially in content synthesis. With the aid of deep learning, co-speech gesture synthesis also bloomed many remarkable methods. These methods attempt to solve this problem by training a conditional model with multi-modal inputs (audio, text, speaker identities).
@@
Co-speech gestures include rhythmic gesture and semantic gesture. Some works attempt to generate semantic gestures in co-speech gesture generation, such as GestureDiffuCLIP [2], LivelySpeaker [52], and Bodyformer [38]. Both GestureDiffuCLIP [2] and LivelySpeaker [52] use CLIP [40] to learn the relationship between semantic text and motion. However, semantic gesture constitutes only a minor portion of the entire sequence and follows a long-tailed distribution, complicating the direct learning of its multi-modal relationships. Moreover, synthesizing gestures that are rhythmically synchronized with the given speech faces a challenge of generalizing to in-the-wild speech scenarios.
@@
Table 1: Comparison of Co-Speech Gesture Datasets. We ignore some datasets with 2D and 3D key-points label. "3D Rot" indicates 3D skeleton with bone rotation. "PGT" means Pseudo Ground Truth.
@@
For addressing the after-mentioned issues, we propose a diffusion-based semantic injection method with large language model for semantic gesture synthesis, which has excellent controllability, interpretability, and generalization ability. It begins with a insight that real-word human conversation contains limited semantic gestures. Besides, for a specific semantic word, such as "wonderful", different people usually present similar gesture such as "opening both arms". During the diffusion reverse process, we can utilize semantic matched gestures as prompts to guide the generative model in synthesizing the appropriate semantic gestures. Therefore, we collect a large-scale semantic motion database, which can include most semantic motion in communication. Large Language Models (LLMs), such as GPT3 [6] and GPT-4, have shown their ability not only as proficient models for natural language processing (NLP) tasks but also as potent instruments for tackling complex problems. Therefore, we introduce large-language model to assign contextually appropriate gestures to the corresponding text. Leveraging the strong semantic analysis capabilities of LLMs, the proposed method is adept at handling various languages.
@@
For a specific speech, different people will present notably different rhythmic gesture. For such many-to-many issue, the best solution is to use deep learning with large-scale dataset to learn their complex relationship. Therefore, it is necessary to build a robust foundational model for rhythmic gesture. Then, the semantic gesture can be injected into the rhythmic gesture with naturalness. However, the co-speech gesture generation research community lack of 3D skeletal data because of expensive motion capture system, which is different from other synthesis tasks, such as text2image, text2speech. Therefore, for improving the basic quality of rhythmic gesture, we collect a large-scale gesture dataset named LSMoG for pre-training that reaches to 400 hours. Table 1 shows the statistical data of co-speech gesture datasets. Some datasets with 2D and 3D key-points label are ignored because they are difficult to use in real applications. To the best of our knowledge, it is the largest dataset in the research of co-speech gesture synthesis field. By the optimization of pre-traing and fine-tuning, the synthesized gesture motion has excellent naturalness and generalization.
The main contributions of our works contains the following aspects:
  1. 1.
@@
    We propose a LLMs-based semantic augmentation method for co-speech gesture synthesis, which can greatly improve the generalization ability in semantic gesture synthesis. For controllable semantic gesture synthesis, we build a set of semantic gesture clips and use LLMs to generate proper candidate gestures for a specific speech content.
  2. 2.
@@
    We build a strong diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels. This dataset is the largest dataset for co-speech gesture synthesis, which contains approximately 400 hours of motion sequences.
  3. 3.
@@
    Extensive experiments show that the proposed method outperforms state-of-the-art methods by a large margin. Especially, the visualization comparison indicates that our method has more stable, expressive, and robust results than other approaches.
2 Related Work
2.1 Co-Speech Gesture Generation
@@
The research of co-speech gesture generation can be divided into two branches, Rule-based methods [9, 35, 10] and learning-based methods [1, 2, 33, 17, 33]. Rule-based works [9, 35, 10] for co-speech gesture synthesis are mainly generated from a pre-defined motion database by keyword matching or other specific rules. These rule-based based methods require lots of human effort in defining gesture units and complex mapping rules, which is costly and inefficient. Besides, the results of rule-based methods are usually lack of smoothness and naturalness.
@@
Co-speech gesture synthesis is a complex problem that requires consideration of the audio, the gesture, and their many-to-many relationships. Benefit from recent advanced developments of deep learning, all of recent methods [1, 2, 33] adopt deep neural networks as a strong tool to learn the complex relationships from audio to gesture in an end-to-end manner. Earlier works [17, 39] explore various network architecture to regress 2D keypoints of the human body movements because of lacking high-quality 3D gesture datasets. These 2D datasets are constructed by using off-the-shelf pose estimator [7] to obtain pseudo 2D gesture annotations from the online videos. However, these 2d results are lack of expressiveness in real application. Recently, some high-quality speech-gesture corpus such as BEAT [32] and ZeroEGGS [16] are released, which contribute to the development of co-speech gesture synthesis. Existing works focus on exploring the effectiveness of different network architectures, including the multi-layer perceptron (MLP) [27], convolutional neural networks (CNN) [20, 48], recurrent neural networks (RNNs) [49, 22, 33, 3] and Transformers [4, 38]. Besides the exploration of model architecture, some approaches [17, 44] dive into researching the connections between co-speech gesture and speech audio, text transcript, speaking style, and speaker identity. Besides, some approaches involve the adversarial training [17, 33], Phase-Guided Motion Matching[45], and reinforcement learning [42] to guarantee realistic results. In addition to gestures, some methods [48, 11, 31] begin to explore the generation of full-body movements, including body gestures and facial expressions.
@@
The rapid development of diffusion models [23, 41] has surpassed the traditional synthesis paradigm such as GAN [19] and VAE [26], showcasing a powerful and realistic characteristic in synthesis. Diffusion-based motion generation also become a popular research direction, such as text2motion [50]. To be specific, text-guided motion can synthesize realistic and expressive results with diffusion model. In co-speech gesture synthesis, recent approaches such as GestureDiffuCLIP[2], LivelySpeaker [52], DiffStyleGesture [44], C2G2[25] are diffusion-based approaches.
2.2 Semantic-aware Co-Speech Gesture Generation
@@
In gesture language, semantic gesture units are crucial for conveying specific intentions, thoughts, emotions, and enhancing gestural expressiveness. For generating more expressive gestures, there are also some attempts such as Gesticulator [27], GestureDiffuCLIP[2], LivelySpeaker[52], Bodyformer [38]. Specifically, GestureDiffuCLIP[2] uses CLIP [21] to align motion and text semantic, then use the text as additional condition to control the semantic gestures. LivelySpeaker[52] is a two-stage framework for semantic and rhythmic gesture generation, which aim at decoupling the semantic gesture generation and rhythmic gesture generation. In semantic gesture generation, they also use a CLIP style module to align the semantic motion and semantic text. Some approaches add semantic text as additional input to fuse semantic into the model, such as Rhythmic Gesticulator [1], Gesticulator [27], FreeTalker [46] and Trimodal [49]. BodyFormer [38] takes both low-level and high-level speech features as input and generates a sequence of realistic 3D body gestures in an auto-regressive manner. The high-level speech features represent the semantic gestures. With the development of LLMs, some approaches also use it to extract motion-related information from textual input, such as GesGPT[15], MoConVQ [47].
3 Method
3.1 Problem Formulation
@@
Given a specific audio, its corresponding text, and identity information, the proposed method can generate high-quality 3D skeletal co-speech gesture results. To be specific, considering a sample of raw audio ${A=\{a_{0},a_{1},...,a_{N}\}}$, we adopt Jukebox [13] to extract its acoustic feature, which is trained on large-scale music datasets. All of the motion sequences =$\mathcal{M}=\{m_{0},m_{1},...,m_{T}\}$ are encoded into discrete tokens by VQVAE[37]. With VQVAE, a motion can be represented by latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{T}\}$. Then, we adopt transformer-based diffusion model as the denoising network to recover the latent embedding by removing the noise. The reverse denoising process of the optimized diffusion model $G$ parameterized by $\theta$ can recover the motion latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{T}\}$ conditioned on the speech audio sequence and identity information. With integration with LLM, it can provide accurate candidate semantic gestures for the speech content. Finally, the candidate semantic gesture embedding are fused into the synthesized motion latent embedding by semantic injection module.
3.2 Semantic Gesture Collection
Through extensive observation of numerous instances, it has been discerned that semantic gestures maintain a universality in gestural communication. Specifically, it is evident that different individuals display similar gestural expressions when articulating identical semantic concepts. This observation provides a feasible methodology for the collection of an extensive array of semantic gesture actions, aimed at covering a comprehensive range of semantic contexts. We collect a large-scale semantic motion clip $G=\{G_{0},G_{1},...,G_{N}\}$ and text $T=\{T_{0},T_{1},...,T_{N}\}$ pairs from BEAT dataset by manual selection. $N$ is the number of the semantic gestures. All of the semantic clip is shorter than 3 seconds and each motion clip represent a whole semantic motion. After selection, we have collected 2,537 semantic motion clips. Some examples of semantic gestures are shown in Figure 2. These motion clips are sorted into a predefined motion semantic system.

### v2

@@ Abstract.
The automated synthesis of high-quality 3D gestures from speech holds significant value for virtual humans and gaming. Previous methods primarily focus on synchronizing gestures with speech rhythm, often neglecting semantic gestures. These semantic gestures are sparse and follow a long-tailed distribution across the gesture sequence, making them challenging to learn in an end-to-end manner. Additionally, generating rhythmically aligned gestures that generalize well to in-the-wild speech remains a significant challenge. To address these issues, we introduce SIGGesture, a novel diffusion-based approach for synthesizing realistic gestures that are both high-quality and semantically pertinent. Specifically, we firstly build a robust diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels. Secondly, we leverage the powerful generalization capabilities of Large Language Models (LLMs) to generate appropriate semantic gestures for various speech transcripts. Finally, we propose a semantic injection module to infuse semantic information into the synthesized results during the diffusion reverse process. Extensive experiments demonstrate that SIGGesture significantly outperforms existing baselines, exhibiting excellent generalization and controllability.
@@
Co-speech gesture synthesis, Semantic gestures, Large Language Models, Diffusion models
journalyear: 2024copyright: acmlicensedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687677isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Animationccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networks
1. Introduction
@@
Body gesture can significantly enhance people’s expressiveness in communication by emphasizing specific intentions or conveying specific semantics (Cassell et al., 1999; Goldin-Meadow and Alibali, 2013; Yoon et al., 2020; Nyatsanga et al., 2023). Additionally, high-quality 3D gesture animation can present vivid virtual characters that help listeners to concentrate and improve the intimacy between humans and virtual characters. Consequently, there is a growing demand for automatic high-quality co-speech gesture synthesis in various applications, including virtual humans, non-player gaming characters, and robot assistants. Many researchers have proposed lots of approaches including rule-based methods (Huang and Mutlu, 2012; Marsella et al., 2013) and data-driven methods (Zhi et al., 2023; Zhu et al., 2023; Ao et al., 2023).
@@
Figure 1. The framework of the proposed method is illustrated in two parts. The upper section details the training processes of diffusion and denoising. The lower part demonstrates the inference process for synthesizing gestures based on the given conditions. Specifically, the audio features and speaker identity are directly fed into the denoising network. Meanwhile, the speech textual content are used as input in the interaction with the LLM, which generates a set of candidate semantic gestures for the subsequent semantic injection process.
@@
3D co-speech gesture generation aims to produce gestures with natural movements, accurate rhythm, and clear semantics based on the given speech. Rule-based methods (Huang and Mutlu, 2012; Marsella et al., 2013) achieve this by using predefined rules, such as associating the keyword ”good” with a simple ”thumbs up” gesture. However, these approaches, lacking naturalness, smoothness, and diversity, can only handle simple semantic gestures. The main reason is that generating co-speech gestures is challenging due to the inherently complex many-to-many mapping. Recently, deep learning with large-scale datasets has achieved significant breakthroughs, especially in content synthesis. With the aid of deep learning, co-speech gesture synthesis has seen the emergence of many remarkable methods. These methods attempt to solve this problem by training a conditional model with multi-modal inputs (audio, text, speaker identities, etc.).
@@
Co-speech gestures include rhythmic gestures and semantic gestures. Some works attempt to generate high-quality co-speech gesture that contains accurate semantic gestures , such as GestureDiffuCLIP (Ao et al., 2023), LivelySpeaker (Zhi et al., 2023), and Bodyformer (Pang et al., 2023). Both GestureDiffuCLIP (Ao et al., 2023) and LivelySpeaker (Zhi et al., 2023) use CLIP (Radford et al., 2021) to learn the relationship between semantic transcripts and motion. However, semantic gestures constitute only a minor portion of the entire sequence and follow a long-tailed distribution, complicating the direct learning of their multi-modal relationships. Moreover, synthesizing gestures rhythmically synchronized with speech faces the challenge of generalizing to in-the-wild scenarios.
@@
Table 1. Comparison of co-speech gesture datasets. We ignore some datasets with 2D and 3D key-points label. ”3D Rot” indicates 3D skeleton with bone rotation. ”PGT” means Pseudo Ground Truth.
@@
To address the aforementioned issues, we propose a diffusion-based semantic injection method with large language model for semantic gesture synthesis, which offers excellent controllability, interpretability, and generalization. It begins with the insight that real-world human conversation contains relatively limited semantic gestures. For example, for specific semantic words like ”wonderful,” different people usually present similar gestures, such as ”opening both arms.” During the diffusion reverse process, we can utilize semantically matched gestures as prompts to guide the generative model in synthesizing appropriate semantic gestures. Therefore, we collect a large-scale semantic motion database that includes most semantic motions in communication. LLMs, such as GPT3 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), have demonstrated their ability not only as proficient models for natural language processing (NLP) tasks but also as potent instruments for tackling complex problems. Therefore, we introduce large-language models to assign contextually appropriate gestures to the corresponding text. Leveraging the strong semantic analysis capabilities of LLMs, the proposed method is adept at handling various languages.
@@
For a specific speech, different people will present notably different rhythmic gestures. For such many-to-many issue, the best solution is to use deep learning with a large-scale dataset to learn their complex relationships. Therefore, it is necessary to build a robust foundational model that can synthesize natural and smooth rhythmic gestures. Then, the semantic gestures can be naturally injected into the rhythmic gestures. However, the co-speech gesture generation research community lacks 3D skeletal data due to the expensive motion capture system, which is different from other synthesis tasks, such as text-to-image, text-to-speech. Therefore, for improving the basic quality of rhythmic gestures, we collected a large-scale gesture dataset named Gesture400 for pre-training, which contains approximately 400 hours of data. Table 1 shows the statistical data of these co-speech gesture datasets. Some datasets with 2D and 3D key-point labels are ignored because they are difficult to use in real applications. To the best of our knowledge, the proposed dataset is the largest dataset in the field of co-speech gesture synthesis research. By the optimization of pre-training and fine-tuning, the synthesized gesture motion exhibits excellent naturalness and generalization.
The main contributions of our work contain the following aspects:
  1. (1)
@@
    We propose an LLMs-based semantic augmentation method for co-speech gesture synthesis, which can greatly improve the generalization ability in semantic gesture synthesis. For controllable semantic gesture synthesis, we build a set of semantic gesture clips and use LLMs to generate appropriate candidate gestures for the specific speech transcripts.
  2. (2)
@@
    We build a robust diffusion-based foundational model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels. This dataset is the largest dataset for co-speech gesture synthesis, containing approximately 400 hours of motion sequences.
  3. (3)
@@
    Extensive experiments show that the proposed method outperforms state-of-the-art methods by a large margin. In particular, the visualization comparisons indicate that our method produces more stable, expressive, and robust results than other approaches.
2. Related Work
2.1. Co-Speech Gesture Generation
@@
The research of co-speech gesture generation can be divided into two branches, rule-based methods (Cassell et al., 1994; Marsella et al., 2013; Cassell et al., 2001) and learning-based methods (Ao et al., 2022, 2023; Liu et al., 2022a; Ginosar et al., 2019). Rule-based works (Cassell et al., 1994; Marsella et al., 2013; Cassell et al., 2001) for co-speech gesture synthesis are mainly generated from a pre-defined motion database by keyword matching or other specific rules. These rule-based methods require lots of human effort in defining gesture units and complex mapping rules, which are costly and inefficient. Besides, the results of rule-based methods are usually lack of smoothness and naturalness.
@@
Co-speech gesture synthesis is a complex problem that requires consideration of the audio, the gesture, and their many-to-many relationships. Benefit from recent advanced developments of deep learning, all of recent methods (Ao et al., 2022, 2023; Liu et al., 2022a) adopt deep neural networks as a strong tool to learn the complex relationships from audio to gesture in an end-to-end manner. Earlier works (Ginosar et al., 2019; Qian et al., 2021) explore various network architecture to regress 2D keypoints of the human body movements because of lacking high-quality 3D gesture datasets. These 2D datasets are constructed by using off-the-shelf pose estimator (Cao et al., 2017) to obtain pseudo 2D gesture annotations from the online videos. However, The 2D results face significant challenges in practical applications.. Recently, some high-quality speech-gesture corpus such as BEAT (Liu et al., 2022b) and ZeroEGGS (Ghorbani et al., 2023) are released, which contribute to the development of co-speech gesture synthesis. Some of existing works focus on exploring the effectiveness of different network architectures, including the Multi-Layer Perceptron (MLP) (Kucherenko et al., 2020), Convolutional Neural Networks (CNN) (Habibie et al., 2021; Yi et al., 2023), Recurrent Neural Networks (RNNs) (Yoon et al., 2020; Hasegawa et al., 2018; Liu et al., 2022a; Bhattacharya et al., 2021a) and Transformers (Bhattacharya et al., 2021b; Pang et al., 2023). Besides the exploration of model architecture, some approaches (Ginosar et al., 2019; Yang et al., 2023b) dive into researching the connections between co-speech gesture and speech audio, text transcript, speaking style, and speaker identity. Furthermore, some approaches involve the adversarial training (Ginosar et al., 2019; Liu et al., 2022a), phase-guided motion matching(Yang et al., 2023a), and reinforcement learning (Sun et al., 2023) to guarantee realistic results. In addition to gestures, some methods (Yi et al., 2023; Chen et al., 2024; Liu et al., 2023) begin to explore the generation of full-body movements, including body gestures and facial expressions.
@@
The rapid development of diffusion models (Ho et al., 2020; Song et al., 2020) has surpassed the traditional synthesis paradigm such as GAN (Goodfellow et al., 2020) and VAE (Kingma and Welling, 2013), showcasing a powerful and realistic characteristic in synthesis. Diffusion-based motion generation also become a popular research direction, such as text2motion (Zhang et al., 2024b). To be specific, text-guided motion can synthesize realistic and expressive results with diffusion model. In co-speech gesture synthesis, recent approaches such as GestureDiffuCLIP(Ao et al., 2023), LivelySpeaker (Zhi et al., 2023), DiffStyleGesture (Yang et al., 2023b), C2G2(Ji et al., 2023), LDA (Alexanderson et al., 2023), Remodiffuse (Zhang et al., 2023a) are diffusion-based approaches.
2.2. Semantic-aware Co-Speech Gesture Generation
@@
In gesture language, semantic gesture units are crucial for conveying specific intentions, thoughts, emotions, and enhancing gestural expressiveness. For generating more expressive gestures, there are also some attempts such as Gesticulator (Kucherenko et al., 2020), GestureDiffuCLIP(Ao et al., 2023), LivelySpeaker(Zhi et al., 2023), Bodyformer (Pang et al., 2023). Specifically, GestureDiffuCLIP(Ao et al., 2023) adopts CLIP (Hafner et al., 2021) to align motion and text semantic, then use the textual feature as additional condition to control the generation process. LivelySpeaker(Zhi et al., 2023) is a two-stage framework for semantic and rhythmic gesture generation, which aim at decoupling the semantic gesture generation and rhythmic gesture generation. In semantic gesture generation, they also use a CLIP style module to align the semantic motion and semantic text. Besides, some approaches add semantic text as additional input to fuse semantic into the model, such as Rhythmic Gesticulator (Ao et al., 2022), Gesticulator (Kucherenko et al., 2020), FreeTalker (Yang et al., 2024) and Trimodal (Yoon et al., 2020). BodyFormer (Pang et al., 2023) takes both low-level and high-level speech features as input and generates a sequence of realistic 3D body gestures in an auto-regressive manner. The high-level speech features represent the semantic gestures. With the development of LLMs, some approaches also use it to extract motion-related information from textual input, such as GesGPT(Gao et al., 2023), MoConVQ (Yao et al., 2023). The proposed method uses Large Language Models (LLMs) to generate proper mapping between the text transcripts and semantic gesture units, which shows excellent performance in both generalization and controllability.
In semantic gesture generation, Semantic Gesticulator(Zhang et al., 2024a) is the most similar work to ours in the same period. Semantic Gesticulator(Zhang et al., 2024a) also attempts to synthesize semantically accurate gestures by introducing a LLM-based semantic gesture units retrieval from a predefined dataset. The gesture generator of Semantic Gesticulator(Zhang et al., 2024a) is based on the GPT-2 (Radford et al., 2019), which predicts the future gesture tokens in an auto-regressive manner. Compared with their work, we use a strong diffusion-based method as fundamental gesture generator to fuse semantic gestures and rhythmic gestures.
3. Method
3.1. Problem Formulation
@@
Given a specific audio, its corresponding text, and identity information, the proposed method can generate high-quality 3D skeletal co-speech gesture results. To be specific, considering a sample of raw audio $A$, we adopt Jukebox (Dhariwal et al., 2020) to extract its acoustic feature, which is pre-trained on large-scale music datasets. All of the motion sequences $\mathcal{M}$ are encoded into discrete tokens by VQVAE (Oord et al., 2017). Details of VQVAE can be found in the appendix. By VQVAE encoding, a motion can be represented by a latent embedding $x_{0}$ Then, we adopt transformer-based diffusion model as the denoising network to recover the latent embedding by removing the noise. The reverse denoising process of the optimized diffusion model $G$ parameterized by $\theta$ can recover the motion latent embedding $x_{0}$ conditioned on the speech audio sequence and identity information. With integration with LLM, it can provide accurate candidate semantic gestures for the speech transcripts. Then, the candidate semantic gesture embedding are fused into the synthesized motion latent embedding by semantic injection module. Finally, the synthesized motion are decoded to continues 3D skeleton rotation by the pre-trained VQVAE.
3.2. Semantic Gesture Collection